# Policymaker meetings as heteroscedasticity shifters:

### v1

@@ 2.1 Motivating example
with $\eta_{t}$ being i.i.d. Gaussian sequence with zero mean and unit variance. The impulse response sequence $C_{h}$ has explicit form $C_{h}=\phi_{1}^{h}$. It is well known that the OLS estimator of $\phi_{1}$ has a non-Gaussian Dickey-Fuller (DF) distribution when $\phi_{1}=1$ (Dickey and Fuller, 1979). Common bootstrap procedures, such as residual bootsrap, are not valid for inference on $\phi_{1}$ and therefore functions of it when $\phi_{1}=1$. However, as proposed in Inoue and Kilian (2002), one can add a time trend or “lag-augment” the model (18),
@@
where the restriction $\phi_{2}=0$ is not imposed. The addition of such an irrelevant regressor results in OLS estimates that have an asymptotic Gaussian marginal distribution of individual components but are not jointly Gaussian. Sims et al. (1990) showed that there exist “forbidden” linear combinations of OLS estimates, corresponding to the so-called generalized cointegrating vectors. Indeed, one can rewrite the model in ADF form
@@
Then, the OLS estimator of $\gamma=(\phi_{1}+\phi_{2})$ will be super-consistent and will have a Dickey-Fuller (DF) asymptotic distribution, while $\hat{\phi}_{2}$ will be asymptotically Gaussian. By the linearity property of the OLS estimators, $\hat{\phi}_{1}=\hat{\gamma}-\hat{\phi}_{2}$. Due to the superconsitency of $\hat{\gamma}$ for 1, the asymptotic distribution of $\hat{\phi}_{1}$ is dominated by the asymptotic Gaussian distribution of $\hat{\phi}_{2}$. As a result, both $\hat{\phi}_{1}$ and $\hat{\phi}_{2}$ have marginal distributions for which the residual bootstrap is valid. Similarly, IRF estimators based on the lag-augmented model (19)
@@
have marginal asymptotic Gaussian distributions with the standard rate of consistency (since asymptotic distribution of $\hat{C}_{2}$ is dominated by the asymptotically Gaussian component $\hat{\phi}_{1}-1$). However the IRF vector is not jointly Gaussian, since the following difference has non-Gaussian distribution
@@
To summarize, the residual bootstrap procedures (without a unit root pre-test and differencing transformation) cannot be directly used for simultaneous inference on IRF in a potential presence of unit root. Similarly, the avaliable proofs for local projection estimators of IRF study asymptotic Gaussian distribution at individual horizons, but do not prove jointly Gaussian distribution (Montiel Olea and Plagborg-Møller, 2021; Xu, 2023). Both approaches require a conservative Bonferroni correction for the construction of simultaneous confidence bounds on multiple horizons (Inoue et al., 2023). [^6] Montiel Olea and Plagborg-Møller (2019) show that sup-t bounds are less conservative than Bonferroni. Other articles concerned with joint inference on IRF in Bayesian VAR framework include Inoue and Kilian (2022); Arias et al. (2023).
In the following sections, we propose an estimation algorithm based on LP, establish consistency and joint asymptotic normality of LP estimators under Assumptions 2-3. Finally, we propose and prove the validity of a novel dependent wild boostrap procedure under the maintained assumptions. This inference procedure provides a way to conduct simultaneous inference on structural IRF for all horizons without pre-tests.
@@ 3.1 Estimation of IRF using LP
Our goal is to estimate
@@
and then compute functions of this parameter.
The first step for our analysis is to estimate $C_{h}$ for $h=1,\dots,H$ and $\Sigma_{\eta}$ using the following hybrid local projection estimator.
@@
    $\hat{\Sigma}_{\eta}=\frac{1}{T-p}\sum_{t=p}^{T-h}\hat{\eta}_{t}\hat{\eta}_{t}^{\prime}.$ (26)
@@
    where $\bar{Z}_{T}=\frac{1}{T}\sum_{t=p}^{T}Z_{t}$.
@@
    For each $h=H_{1}+1,\dots,H_{2}$ compute recursively
    $\hat{C}^{BR}_{h}=\hat{C}^{BR}_{h}(\hat{\theta}(H_{1}))=\sum_{\ell=1}^{i}\hat{A}^{BR}_{\ell}\hat{C}^{*}_{i-\ell},$ (30)
@@
Algorithm 1 gives estimates of $C_{h}$ for the first $H_{1}$ horizons that are identical to the local projection estimator of Jordà (2005) as Theorem 2 below shows. Since we are interested in simultaneous inference on all relevant parameters of an SVAR, this algorithm imposes the VAR(p) structure on impulse responses of longer horizons. In doing so, we reduce the problem of inference about IRF at long horizons to the problem of simultaneous inference on a finite set of parameters $C_{1},\dots,C_{H_{1}}$. This is a much easier problem and can be addressed by our bootstrap procedure, which will be outlined later.
Note that, in general, the backward recursion estimators of $A_{i}$ are different from the corresponding LS estimators. The LS estimates may be super-consitent for some linear combinations of $A$, but it comes at the expense of non-standard asymptotic distributions that are hard to use for confidence sets (Mikusheva, 2012). On the contrary, $\hat{A}^{BR}_{i}$ are polynomial functions of asymptotically Gaussian estimators $\hat{C}_{h}$. Therefore, a bootstrap procedure can be used to estimate the bias and variance of $\hat{A}^{BR}_{i}$. This approach allows us to bypass any need to pre-test data for presence of deterministic trends, unit roots, and cointegration relations and avoid unnecessary data transformations (which may reduce efficiency and introduce pretesting biases).
@@
Suppose Assumptions 2 and 3 hold. Then local projection estimators of $C_{h}$ using regressions
$y_{t+h}=C_{h}y_{t}+\tilde{V}(h)\mu_{t-1}+\sum_{i=1}^{p}\tilde{A}_{i}(h)y_{t-i}+\xi_{t+h}(h),$ (31)
@@
$\hat{C}_{h}^{\prime}=(\sum_{t=p}^{T-h}\hat{\eta}_{t}\hat{\eta}_{t}^{\prime})^{-1}\sum_{t=p}^{T-h}\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+h}(h)$ (32)
where $\hat{\eta}_{t}=\hat{\eta}_{t}(0)$ and $\hat{\eta}_{t+h}(h)$ are regression residuals from $h$-step ahead autoregressions defined in (25). Moreover, $\hat{C}_{h}\stackrel{{\scriptstyle p}}{{\rightarrow}}C_{h}$ for all $h$ and $\hat{\Sigma}_{\eta}\stackrel{{\scriptstyle p}}{{\rightarrow}}\Sigma_{\eta}$.
@@
  3. Step 3.
    Compute forecast error variance decomposition for each $r=1,\dots,n$ and $H$ using
    $\widehat{FEVD}_{r}(H)=\frac{\sum_{h=0}^{H}(e^{\prime}_{r}\hat{\psi}_{h})^{2}}{\sum_{h=0}^{H}e^{\prime}_{r}\hat{C}_{h}\hat{\Sigma}_{\eta}\hat{C}^{\prime}_{h}\iota}.$ (34)
Algorithm 2 is consistent for estimation of structural IRF and FEVD for the shock of interest as follows from
@@ 3.2 HAC-inference
In the previous section, we obtained the asymptotic Gaussian distribution and the corresponding covariance matrix for $\hat{\theta}(H_{1})$. This covariance matrix can be estimated using heteroscedasticity and autocorrealtion consistent (HAC) covariance estimator as the following theorem shows.
@@
with $\Omega(H)\triangleq\sum^{\infty}_{\ell=-\infty}\mathbb{E}\Xi_{t}(H)\Xi^{\prime}_{t+\ell}(H)$, where
$\Xi_{t}(H)\triangleq\begin{pmatrix}vec(\eta_{t}\eta_{t}^{\prime}-\Sigma_{\eta})\\ vec(\Sigma_{\eta}^{-1}\eta_{t}\eta_{t+1}^{\prime})\\ \cdots\\ vec(\Sigma_{\eta}^{-1}\sum_{s=0}^{H-1}\eta_{t}\eta_{t+H-s}^{\prime}C_{s}^{\prime})\\ (\eta_{t}\eta_{1t}-\mathbb{E}\eta_{t}\eta_{1t})(Z_{t}-\mathbb{E}Z_{t})-\gamma\end{pmatrix}.$ (36)
@@
The variance-covariance matrix $\Omega$ can be estimated using HAC estimator along the lines in Andrews (1991). We introduce
@@
Even though we had to modify the original proofs of Andrews (1991) to account for non-stationarity of $y_{t}$, the asymptotic theory for optimal selection of kernels and bandwidths still holds. This is due to the fact that the infeasible analog of $\hat{\Omega}(H)$ that uses the true $H$-step innovations $\eta_{t}(H)$ has the same properties regardless of the presence of unit roots and deterministic trends in the VAR model. In particular, one can use the conventional rule $B_{T}=\{0.75\sqrt[3]{T}$ that is appropriate for Bartlett kernel. 777 Recent literature emphasized the fact that when data is very persistent, HAC estimators do not perform well (see, fore example Lazarus et al., 2021). In particular, in the near-to-unity asymptotics, one should use a very large bandwidth $B_{T}$ to control the size of $t$ tests. However, if the innovations are two-sided MDS, then Montiel Olea and Plagborg-Møller (2021) showed that one can use $B_{T}=0$ for inference on individual coefficients $\hat{C}_{h}$. Even under these stronger assumptions, joint inference on the entire vector $\theta(H_{1})$ still requires $B_{T}$ to grow with the sample size to account for potential presence of conditional hetersedasticity of unknown form.
@@ 3.3 Dependent wild bootstrap
Using the Delta method, an asymptotic normal distribution can be derived for other estimators, including $\hat{A}^{BR}_{j}$, $\hat{C}^{BR}_{h}$ at long horizons $h$, and $\widehat{FEVD}(h)$. The wild bootstrap approach of Shao (2010) can be used as a convincing numerical implementation of the delta method. A particular way to perform bootstrap is given in the following algorithm.
@@
    Select bandwidth $B_{T}$. For example, $B_{T}=0.75\sqrt[3]{T}$
@@
    Generate i.i.d. $\zeta^{s}_{i}\sim N(0,\frac{1}{B_{T}})$ for $i=-B_{T}+1,-B_{T}+2,...,0,1,...T$
@@
    Recursively compute $u^{s}_{t}=\zeta^{s}_{t}+\zeta^{s}_{t-1}+...+\zeta^{s}_{t-B_{T}+1}$ , $t=1,...T$
@@
    Compute $\hat{A}^{BR,s}_{i}=A^{BR}_{i}(\hat{\theta}^{s}(H_{1}))$, $\hat{C}^{BR,s}_{h}=C^{BR}_{h}(\hat{\theta}^{s}(H_{1}))$, $\hat{b}^{s}_{\cdot 1}(\hat{\theta}^{s}(H_{1}))$ using Algorithms 1 and 2.
@@
Suppose Assumption 1-3 holds and $B_{T}^{2}/T\to 0$. Then
@@
To our knowledge, this is the first joint bootstrap resampling procedure for LP estimators. The validity of the dependent wild bootstrap is based on the fact that it mimics the asymptotic Gaussian distribution of $\hat{\theta}(H_{1})$ with a covariance matrix that coincides with the HAC estimator given in Theorem 5. This argument does not impose any parametric model on $\eta_{t}$ and $Z_{t}$ and is generally applicable under the maintained weak dependence assumption (Assumption 2.2). Note that the algorithm uses $\hat{\Xi}_{t}$ that depend on possibly non-stationary processes with strong dependence $Y_{t}$. As a result, conventional block bootstrap procedures are not directly applicable to $\hat{\Xi}_{t}$. Unlike the popular recursive residual-based bootstrap procedures of Gonçalves and Kilian (2004), we do not assume stability (or stationarity) of $Y_{t}$ and allow for asymmetric conditional heteroscedasticity in $\eta_{t}$.

### v2

@@ 2.1 Motivating example
with $\eta_{t}$ being i.i.d. Gaussian sequence with zero mean and unit variance. The impulse response sequence $C_{h}$ has explicit form $C_{h}=\phi_{1}^{h}$. It is well known that the OLS estimator of $\phi_{1}$ has a non-Gaussian Dickey-Fuller (DF) distribution when $\phi_{1}=1$ (Dickey and Fuller, 1979). The common bootstrap procedures, such as residual bootstrap, are not valid for inference on $\phi_{1}$ and therefore functions of it when $\phi_{1}=1$. However, as proposed in Inoue and Kilian (2002), one can add a time trend or “lag-augment” the model (18),
@@
where the restriction $\phi_{2}=0$ is not imposed. The addition of such irrelevant regressor results in OLS estimates that have an asymptotic Gaussian marginal distribution of individual components but are not jointly Gaussian. Sims et al. (1990) showed that there exist “forbidden” linear combinations of OLS estimates, corresponding to the so-called generalized cointegrating vectors. Indeed, one can rewrite the model in ADF form
@@
Then, the OLS estimator of the corresponding population parameter $\gamma=(\phi_{1}+\phi_{2})$ will be super-consistent and will have a Dickey-Fuller (DF) asymptotic distribution, while $\hat{\phi}_{2}$ will be asymptotically Gaussian. By the linearity property of the OLS estimators, $\hat{\phi}_{1}=\hat{\gamma}-\hat{\phi}_{2}$. Due to the super-consistency of $\hat{\gamma}$ for 1, the asymptotic distribution of $\hat{\phi}_{1}$ is dominated by the asymptotic Gaussian distribution of $\hat{\phi}_{2}$. As a result, both $\hat{\phi}_{1}$ and $\hat{\phi}_{2}$ have marginal distributions for which the residual bootstrap is valid. Similarly, IRF estimators based on the lag-augmented model (19)
@@
have marginal asymptotic Gaussian distributions with the standard rate of consistency (since the asymptotic distribution of $\hat{C}_{2}$ is dominated by the asymptotically Gaussian component $\hat{\phi}_{1}-1$). However, the IRF vector is not jointly Gaussian, since the following difference has a non-Gaussian distribution
@@
To summarize, the residual bootstrap procedures (without a unit root pre-test and differencing transformation) cannot be directly used for simultaneous inference on IRF in the potential presence of a unit root. Similarly, the available proofs for local projection estimators of IRF study asymptotic Gaussian distribution at individual horizons, but do not prove jointly Gaussian distribution (Montiel Olea and Plagborg-Møller, 2021; Xu, 2023). Both approaches require a conservative Bonferroni correction for the construction of simultaneous confidence bounds on multiple horizons (Inoue et al., 2023). [^8] Montiel Olea and Plagborg-Møller (2019) show that sup-t bounds are less conservative than Bonferroni. Other articles concerned with joint inference on IRF in the Bayesian VAR framework include Inoue and Kilian (2022); Arias et al. (2023).
In the following sections, we propose an estimation algorithm based on LP, and establish consistency and joint asymptotic normality of LP estimators under Assumptions 2-3. Finally, we propose and prove the validity of a novel dependent wild bootstrap procedure under the maintained assumptions. This inference procedure provides a way to conduct simultaneous inference on structural IRF for all horizons without pre-tests.
@@ 3.1 Estimation of IRF using LP
Our goal is to estimate the vector of parameters
@@
and then compute the functions of this parameter.
The first step for our analysis is to estimate $C_{h}$ for $h=1,\dots,H$, and $\Sigma_{\eta}$ using the following hybrid local projection estimator.
@@
    $\hat{\Sigma}_{\eta}=\frac{1}{T-p}\sum_{t=p}^{T}\hat{\eta}_{t}\hat{\eta}_{t}^{\prime}.$ (26)
@@
    where $\bar{Z}_{T}\triangleq\frac{1}{T-p}\sum_{t=p}^{T}Z_{t}$ and $\overline{\hat{\eta}_{t}\hat{\eta}_{1t}}\triangleq\frac{1}{T-p}\sum_{t=p}^{T}\hat{\eta}_{t}\hat{\eta}_{1t}$.
@@
    For each $h=H_{1}+1,\dots,H_{2}$ compute recursively as follows
    $\hat{C}^{BR}_{h}=\hat{C}^{BR}_{h}(\hat{\theta}(H_{1}))=\sum_{\ell=1}^{h}\hat{A}^{BR}_{\ell}\hat{C}^{*}_{h-\ell},$ (30)
@@
Algorithm 1 gives estimates of $C_{h}$ for the first $H_{1}$ horizons. These estimates are identical to the local projection estimator of Jordà (2005) as Theorem 2 below shows. Since we are interested in simultaneous inference on all relevant parameters of an SVAR, this algorithm imposes the VAR(p) structure on impulse responses of longer horizons, $H_{1}\leq h\leq H_{2}$. In doing so, we reduce the problem of inference about IRF at long horizons to the problem of simultaneous inference on a finite set of parameters $C_{1},\dots,C_{H_{1}}$. This is a much easier problem, and our bootstrap procedure, which will be outlined later, can address it.
Note that, in general, the backward recursion estimators of $A_{i}$ are different from the corresponding LS estimators. The LS estimates may be super-consistent for some linear combinations of $A$, but come at the expense of non-standard asymptotic distributions that are hard to use for confidence sets (Mikusheva, 2012). In contrast, $\hat{A}^{BR}_{i}$ are polynomial functions of asymptotically Gaussian estimators $\hat{C}_{h}$. Therefore, a bootstrap procedure can estimate the bias and variance of $\hat{A}^{BR}_{i}$. This approach allows us to bypass any need to pre-test data for the presence of deterministic trends, unit roots, and cointegration relations and avoid unnecessary data transformations. Such pretests may introduce pretesting biases and reduce efficiency of the resulting estimators if the transformations turn out to be redundant.
@@
Suppose Assumptions 2 and 3 hold. Then the local projection estimators of $C_{h}$ using regressions
$y_{t+h}=C_{h}y_{t}+\tilde{V}(h)\mu_{t-1}+\sum_{i=1}^{p}\tilde{A}_{i}(h)y_{t-i}+\xi_{t+h}(h)$ (31)
@@
$\hat{C}_{h}^{\prime}=(\sum_{t=p}^{T-h}\hat{\eta}_{t}\hat{\eta}_{t}^{\prime})^{-1}\sum_{t=p}^{T-h}\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+h}(h),$ (32)
where $\hat{\eta}_{t}\triangleq\hat{\eta}_{t}(0)$ and $\hat{\eta}_{t+h}(h)$ are regression residuals from $h$-step ahead autoregressions defined in (25). Moreover, $\hat{C}_{h}\stackrel{{\scriptstyle p}}{{\rightarrow}}C_{h}$ for all $h$ and $\hat{\Sigma}_{\eta}\stackrel{{\scriptstyle p}}{{\rightarrow}}\Sigma_{\eta}$.
@@
Algorithm 2 is consistent for the estimation of structural IRF for the shock of interest as follows from
@@
Remark 1.
One can also compute forecast error variance decomposition for each $r=1,\dots,n$ and $H$ using formula
$\widehat{FEVD}_{r}(H)=\frac{\sum_{h=0}^{H-1}(e^{\prime}_{r}\hat{\psi}_{h})^{2}}{\sum_{h=0}^{H}e^{\prime}_{r}\hat{C}_{h}\hat{\Sigma}_{\eta}\hat{C}^{\prime}_{h}\iota},$ (34)
where $\iota$ is a vector of ones.
@@ 3.2 HAC-inference
In the previous section, we obtained the asymptotic Gaussian distribution and the corresponding covariance matrix for $\hat{\theta}(H_{1})$. This covariance matrix can be estimated using heteroscedasticity and autocorrelation consistent (HAC) covariance estimator as the following theorem shows.
@@
with $\Omega(H_{1})\triangleq\sum^{\infty}_{\ell=-\infty}\mathbb{E}\Xi_{t}(H_{1})\Xi^{\prime}_{t+\ell}(H_{1})$, where
$\Xi_{t}(H_{1})\triangleq\begin{pmatrix}vec(\eta_{t}\eta_{t}^{\prime}-\Sigma_{\eta})\\ vec(\Sigma_{\eta}^{-1}\eta_{t}\eta_{t+1}^{\prime})\\ \cdots\\ vec(\Sigma_{\eta}^{-1}\sum_{s=0}^{H_{1}-1}\eta_{t}\eta_{t+H_{1}-s}^{\prime}C_{s}^{\prime})\\ (\eta_{t}\eta_{1t}-\mathbb{E}\eta_{t}\eta_{1t})(Z_{t}-\mathbb{E}Z_{t})-\gamma\end{pmatrix},$ (36)
and $\eta_{1t}$ is the first element of $\eta_{t}$.
@@
The variance-covariance matrix $\Omega$ can be estimated using the HAC estimator along the lines in Andrews (1991). We introduce
@@
Even though we had to modify the original proofs of Andrews (1991) to account for the nonstationarity of $y_{t}$, the asymptotic theory for optimal selection of kernels and bandwidths still holds. This is because the infeasible analog of $\hat{\Omega}(H)$ that uses the true $H$-step innovations $\eta_{t}(H)$ has the same properties regardless of the presence of unit roots and deterministic trends in the VAR model. In particular, one can use the conventional rule $B_{T}=\{0.75\sqrt[3]{T}\}$ that is appropriate for the Bartlett kernel.
Recent literature emphasized the fact that when data is very persistent, HAC estimators do not perform well (see, for example Lazarus et al., 2021). In particular, in the near-to-unity asymptotics, one should use a very large bandwidth $B_{T}$ to control the size of $t$ tests. This is not the case under our Assumption 2.2 (weak dependence) for the innovations $\eta_{t}$. If, in addition, the innovations are two-sided MDS, then Montiel Olea and Plagborg-Møller (2021) showed that one can use $B_{T}=0$ for inference on individual coefficients $\hat{C}_{h}$. However, even under these stronger assumptions, the joint inference on the entire vector $\theta(H_{1})$ still requires $B_{T}$ to grow with the sample size to account for the potential presence of conditional heteroscedasticity of unknown form.
@@ 3.3 Dependent wild bootstrap
Using the Delta method, an asymptotic normal distribution can be derived for other estimators, including $\hat{A}^{BR}_{j}$ and $\hat{C}^{BR}_{h}$ at long horizons $h$999 Another function of interest is the FEVD. Gorodnichenko and Lee (2020) studied inference on FEVD using local projections and stationary VAR-bootstrap under assumption of homoscedasticity of shocks. They show that block bootstrap of Kilian and Kim (2011) can also be used for inference on FEVD under the assumption of stationary data. The dependent wild bootstrap approach can also be applied to and $\widehat{FEVD}(h)$ without requiring stationarity of the VAR model. The dependent wild bootstrap approach of Shao (2010) can be used as a convincing numerical implementation of the delta method. A particular way to perform bootstrap is given in the following algorithm.
@@
    Select bandwidth $B_{T}$ according to Theorem 5. For example, $B_{T}=0.75\sqrt[3]{T}$.
@@
    Generate i.i.d. $\zeta^{s}_{t}\sim N(0,\frac{1}{B_{T}})$ for $t=-B_{T}+1,-B_{T}+2,...,0,1,...T$.
@@
    Recursively compute $u^{s}_{t}=\zeta^{s}_{t}+\zeta^{s}_{t-1}+...+\zeta^{s}_{t-B_{T}+1}$ , $t=1,...T$.
@@
    Compute $\hat{A}^{BR,s}_{i}=A^{BR}_{i}(\hat{\theta}^{s}(H_{1}))$, $\hat{C}^{BR,s}_{h}=C^{BR}_{h}(\hat{\theta}^{s}(H_{1}))$, $\hat{b}^{s}_{\cdot 1}(\hat{\theta}^{s}(H_{1}))$ for $h\geq H_{1}$ using Algorithms 1 and 2.
@@
Suppose Assumption 1-3 hold and $B_{T}^{2}/T\to 0$. Then
@@
To our knowledge, this is the first joint bootstrap resampling procedure for LP estimators. The validity of the dependent wild bootstrap is based on the fact that it mimics the asymptotic Gaussian distribution of $\hat{\theta}(H_{1})$ with a covariance matrix that coincides with the HAC estimator given in Theorem 5. This argument does not impose any parametric model on $\eta_{t}$ and $Z_{t}$ and is generally applicable under the maintained Assumption 2.2 (weak dependence of innovations). Note that the algorithm uses $\hat{\Xi}_{t}$ that depend on possibly nonstationary processes with strong dependence, $Y_{t}$. As a result, conventional block bootstrap procedures are not directly applicable to $\hat{\Xi}_{t}$. Unlike the popular recursive residual-based bootstrap procedures of Gonçalves and Kilian (2004), we do not assume stability (or stationarity) of $Y_{t}$ and allow for asymmetric conditional heteroscedasticity in $\eta_{t}$.